OpenStudy (anonymous):

What statement correctly describes the key features of the graph of f(x) = 4(one half)x + 1 − 3? a.Y-intercept of (0, −1), starts down on the left, gets closer to y = −3 on the right b.Y-intercept of (0, −1), starts up on the left, c.gets closer to y = −3 on the right Y-intercept of (0, 1), starts up on the left, gets closer to y = −3 on the right d.Y-intercept of (0, 1), starts down on the left, gets closer to y = −3 on the right

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@KrissieG123

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

@quickstudent

OpenStudy (anonymous):

y=mx+b, m = slope, b = y - intercept (0, something)

OpenStudy (anonymous):

i know the y-intercept x=0 , when i plug in x=0 in the equation i get the answer -1 so it's (0, -1) correct? @iambatman

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@Luigi0210

OpenStudy (anonymous):

\[f(x) = 4 \frac{ 1 }{ 2 } x+1-3\] this?

OpenStudy (anonymous):

y=f(x)

OpenStudy (anonymous):

x is raised to the power , i solved it like this 4(1/2)^0+1 -3 = -1

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@MTALHAHASSAN2

OpenStudy (anonymous):

@mathrulezz @mathmath333

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{is this the equation }\hspace{.33em}\\~\\ f(x) = 4\cfrac{1}{2}x + 1 − 3 \hspace{.33em}\\~\\ \normalsize \text{or this }\hspace{.33em}\\~\\ f(x) = \dfrac{4x}{2} + 1 − 3 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

this is the equation : f(x) = 4(1/2)^x+1 -3

OpenStudy (anonymous):

so the y-intercept is x=0 when i plug that in i get f(x) which is y= -1 , and the answer is i think (0,-1) correct?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{u mean this one }\hspace{.33em}\\~\\ f(x) = 4\times \left(\dfrac{1}{2}\right)^{x + 1 }− 3 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

exactly and -3 after ^x+1

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@Whitemonsterbunny17

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{first check the y -intercept,put x=0 }\hspace{.33em}\\~\\ f(x) = 4\dfrac{1}{2}^{x + 1 }− 3 \hspace{.33em}\\~\\ y= 4\dfrac{1}{2}^{x + 1 }− 3 \hspace{.33em}\\~\\ y= 4\dfrac{1}{2}^{0+ 1 }− 3 \hspace{.33em}\\~\\ y= 4\dfrac{1}{2}− 3 \hspace{.33em}\\~\\ y= 2− 3 =-1\hspace{.33em}\\~\\ \normalsize \text{so y -intercept=(0,-1) }\hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

yup!! , and that's what i exactly got but there is two options that has the same answer and i'm confused

OpenStudy (mathmath333):

let's see...

OpenStudy (mathmath333):

we want to see if its starts on left or right

OpenStudy (mathmath333):

i mean starts down or up in left

OpenStudy (mathmath333):

lol i m confused

OpenStudy (anonymous):

i have graphed it i'll show you , maybe you can help me even tho i think it stats up

OpenStudy (mathmath333):

https://www.desmos.com/calculator/eg5xnoahsq

OpenStudy (mathmath333):

here i think u have to put larger values of positive ans negative values f(x) that is y and test.for some higher negative values of y u will get x undefined.

OpenStudy (mathmath333):

so its option b.) if we graph it

OpenStudy (anonymous):

yeah and i was thinking the same thing too because it will never start from the bottom anyway and the values i have plugged got me positive number and for x if we plug negative numbers we can't raise a number to the negative anyway @mathmath333

OpenStudy (anonymous):

Thank you so much the answer b.Y-intercept of (0, −1), starts up on the left, gets closer to y = −3 on the right was correct and i just got a 100%

OpenStudy (mathmath333):

see the difference between positive and negative https://www.desmos.com/calculator/aj0ariaifg

OpenStudy (anonymous):

yeah i see but for this problem it started from the top , anyways thanks a lot and you have helped me the last time and you're so helpful and such an awesome person man!

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

yw